Supposed five coins are tossed.let x be the random variable representing the number of head occur. Find the probability of each values of the random variable.

1
Expert's answer
2022-01-31T18:05:28-0500

Solution:

We will assume that the probability of getting heads and tails is the same: "p = q = \frac{1}{2}".

Using Bernoulli's formula, we find the probability that 0, 1, 2, 3, 4, and 5 heads will land:

"P(0) = {q^5} = {\left( {\frac{1}{2}} \right)^5} = \frac{1}{{32}}"


"P(1) = C_5^1p{q^4} = 5 \cdot {\left( {\frac{1}{2}} \right)^5} = \frac{5}{{32}}"


"P(2) = C_5^2{p^2}{q^3} = 10 \cdot {\left( {\frac{1}{2}} \right)^5} = \frac{{10}}{{32}}"


"P(3) = C_5^3{p^3}{q^2} = 10 \cdot {\left( {\frac{1}{2}} \right)^5} = \frac{{10}}{{32}}"


"P(4) = C_5^4{p^4}q = 5 \cdot {\left( {\frac{1}{2}} \right)^5} = \frac{5}{{32}}"


"P(5) = {p^5} = {\left( {\frac{1}{2}} \right)^5} = \frac{1}{{32}}"


We get the distribution law

"\begin{matrix}
X&0&1&2&3&4&5\\
p&{\frac{1}{{32}}}&{\frac{5}{{32}}}&{\frac{{10}}{{32}}}&{\frac{{10}}{{32}}}&{\frac{5}{{32}}}&{\frac{1}{{32}}}
\end{matrix}"